OpenStudy (credmond):

A. Find the critical numbers B. the largest open intervals where the function is increasing C. the largest open intervals where it is decreasing f(x)= 2/3x^3 - x^2 - 4x + 2

OpenStudy (anonymous):

is it \[f(x)=\frac{2}{3}x^3-x^2-4x+2\]?

OpenStudy (credmond):

Yep!

OpenStudy (anonymous):

k did you take the derivative?

OpenStudy (credmond):

No, I didn't really know where to start :/

OpenStudy (anonymous):

take the derivatives using the power rule

OpenStudy (anonymous):

clear or no?

OpenStudy (credmond):

Yeah so the derivative would be 2x^2- 2x -4 right?

OpenStudy (anonymous):

right now set it equal to zero and solve

OpenStudy (anonymous):

you got that part? this problem has been cooked up, so the derivative factors easily

OpenStudy (credmond):

so I got x = 2, and -2 right

OpenStudy (anonymous):

no

OpenStudy (credmond):

oh wait... 2 and -1

OpenStudy (anonymous):

\[2x^2- 2x -4 =2(x^2-x-2)=2(x+1)(x-2)\] \[(x+1)(x-2)=0\] so ... yeah, those

OpenStudy (credmond):

haha okay!

OpenStudy (anonymous):

ok the critical numbers at \(-1\) and \(2\)

OpenStudy (anonymous):

you don't really need anything else to finish the problem you know the two critical numbers, and you know that a cubic polynomial with positive leading coefficient looks something like ths 

OpenStudy (anonymous):

i.e. it is increasing, then decreasing, then increasing you could also check the sign of the derivative on the open intervals \((-\infty, -1),(-1,2),(2,\infty)\) to see where it is positive and where it is negative you good from there, or you need more help?

OpenStudy (credmond):

Okay, cool. I think I'm good from here!

OpenStudy (anonymous):

k good luck

OpenStudy (credmond):

Thank you so much for the help! I really appreciate it! :)

OpenStudy (anonymous):

yw